# Recent Advances in Speech Language Models: A Survey

### v1

@@ Abstract
Large Language Models (LLMs) have recently garnered significant attention, primarily for their capabilities in text-based interactions. However, natural human interaction often relies on speech, necessitating a shift towards voice-based models. A straightforward approach to achieve this involves a pipeline of “Automatic Speech Recognition (ASR) + LLM + Text-to-Speech (TTS)”, where input speech is transcribed to text, processed by an LLM, and then converted back to speech. Despite being straightforward, this method suffers from inherent limitations, such as information loss during modality conversion and error accumulation across the three stages. To address these issues, Speech Language Models (SpeechLMs)—end-to-end models that generate speech without converting from text—have emerged as a promising alternative. This survey paper provides the first comprehensive overview of recent methodologies for constructing SpeechLMs, detailing the key components of their architecture and the various training recipes integral to their development. Additionally, we systematically survey the various capabilities of SpeechLMs, categorize the evaluation metrics for SpeechLMs, and discuss the challenges and future research directions in this rapidly evolving field.
1 Introduction
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities in generating text and performing a wide array of natural language processing tasks Achiam et al. (2023); Dubey et al. (2024a); Zhang et al. (2022b), serving as powerful foundation models for AI-driven language understanding and generation. Their success has also spurred numerous applications in various other domains, yet the reliance solely on text-based modalities presents a significant limitation. This leads to the development of speech-based generative models, which allow to interact with humans more naturally and intuitively. The inclusion of speech not only facilitates real-time voice interactions but also enriches communication by combining both text and speech information Nguyen et al. (2023b, 2024).
@@
Given the extensive mutual information between text and speech, it is natural to modify existing LLMs to enable speech interaction capabilities. A straightforward approach is to adopt an “Automatic Speech Recognition (ASR) + LLM + Text-to-Speech (TTS)” framework (Figure 1a) Huang et al. (2024). In this setup, the user’s spoken input is first processed by the ASR module, which converts it into text. The LLM then generates a text response based on this transcription. Finally, the TTS module transforms the text response back into speech, which is played back to the user. However, this naive solution mainly suffers from the following two problems. 1) Information loss. Speech signals not only contain semantic information (i.e., the meaning of the speech) but also paralinguistic information (e.g., pitch, timbre, tonality, etc.). Putting a text-only LLM in the middle will cause the complete loss of paralinguistic information in the input speech Zhang et al. (2023a). 2) Cumulative error. A staged approach like this can easily lead to cumulative errors throughout the pipeline, particularly in the ASR-LLM stage Fathullah et al. (2024). Specifically, transcription errors that occur when converting speech to text in the ASR module can negatively impact the language generation performance of the LLM.
@@
The limitations of the naive ASR + LLM + TTS framework have led to the development of Speech Language Models (SpeechLMs, Figure 1b). Unlike the naive framework, SpeechLMs directly encode speech waveforms into discrete tokens, capturing essential features and information from audio (section 3.1). Although individual speech tokens may not carry word-level semantic meaning, they capture the semantic information of speech utterances and retain valuable paralinguistic information, which prevents the information loss. SpeechLMs then model these tokens autoregressively, without solely relying on text input, which allows them to use the additional paralinguistic information to generate more expressive and nuanced speech (section 3.2). Finally, the generated tokens are synthesized back to speech (section 3.3). By working directly with the encoded speech tokens, SpeechLMs effectively mitigate the cumulative errors, as their training is integrated with the speech encoding, whereas the training of LLMs (language modeling) is completely independent of the ASR (speech recognition) module in the naive framework.
@@
Beyond basic conversational abilities, SpeechLMs hold the potential to undertake more complex tasks, such as encoding speaker-specific information and emotional nuances (Figure 2). This capability allows SpeechLMs to distinguish between different speakers during a conversation and to comprehend and generate speech imbued with specific emotional tones. Such advancements are crucial for applications in areas like personalized assistants, emotion-aware systems, and more nuanced human-computer interaction scenarios. Furthermore, SpeechLMs can be designed to enable real-time voice interaction, where the model can be interrupted by humans or choose to speak while the user is still speaking, which resembles the pattern of human conversations more closely.
@@
In this survey, we present the first comprehensive overview of recent endeavors in constructing SpeechLMs. We explore the various components that constitute their architecture (section 3) and the training recipes (table 2) involved in their development. we aim to elucidate the current state of the field by analyzing these models from the above perspectives. Additionally, we survey the downstream applications of SpeechLMs (section 5), classify metrics to evaluate SpeechLMs (section 6), discuss the challenges encountered in this rapidly evolving area, and outline promising future research directions that could drive further advancements in SpeechLM technology (section 7). Our contributions are summarized as follows:
@@ 2 Problem Formulation
In this section, we provide a formal definition of Speech Language Models. A Speech Language Model (SpeechLM) is an autoregressive foundation model that processes and generates speech data, utilizing contextual understanding for coherent sequence generation. It supports both speech and text modalities, such as speech-in-text-out, text-in-speech-out, or speech-in-speech-out, enabling a wide range of tasks with context-aware capabilities. We note that the concept of SpeechLM is in contrast to traditional text-based language models, such as LLM, where the only modality being processed within the model is text. Therefore, to avoid confusion, we call those text-based language models TextLMs throughout this survey.
We offer a unified framework in which SpeechLMs can process and generate speech data, text data, or even interleaved speech and text data. Specifically, an audio waveform $\textbf{a}=(a_{1},a_{2},\ldots,a_{Q})$ consists of a sequence of audio samples $a_{i}\in\mathbb{R}$ of length $Q$, where $1\leq q\leq Q$. Similarly, a text span $\textbf{t}=(t_{1},t_{2},\ldots,t_{K})$ consists of a sequence of text tokens $t_{j}$ (word, subword, character, etc.) of length $K$. Let $\textbf{M}=(M_{1},M_{2},\ldots,M_{N})$ denote a multimodal sequence of length $N$, where each element $M_{i}\in\{a_{i},t_{j}\}$. We define $\textbf{M}^{\text{in}}=(M_{1}^{\text{in}},M_{2}^{\text{in}},\ldots,M_{N_{\text{in}}}^{\text{in}})$ as the input multimodal sequence and $\textbf{M}^{\text{out}}=(M_{1}^{\text{out}},M_{2}^{\text{out}},\ldots,M_{N_{\text{out}}}^{\text{out}})$ as the output multimodal sequence, where $N_{\text{in}}\geq 0$ and $N_{\text{out}}\geq 0$. Then, A SpeechLM parameterized by $\theta$ can then be represented as:
@@
3 Components in SpeechLM
There are three main components within a SpeechLM, namely speech tokenizer, language model, and token-to-speech synthesizer (vocoder), as illustrated in Figure 1. The fundamental reason for such a three-staged design pattern is to use the language modeling architecture (e.g., decoder-only transformer) to model speech autoregressively in the format of audio waveforms. Since both the input and output of a language model are discrete tokens, additional modules need to be attached to the language model to handle the I/O format. Specifically, the speech tokenizer first transforms continuous audio waveforms into discrete tokens to serve as input to the language model, then the language model performs the next-token prediction based on the input speech tokens. Finally, the vocoder transforms the discrete tokens outputted by the language model back into audio waveforms. We note that our focus here is on how the three components are grouped together to form a SpeechLM rather than a comprehensive overview of each component. Therefore, for speech tokenizer and vocoder, we mainly summarize the methods used in existing SpeechLMs. Table 1 summarizes the popular choices of the three components in various SpeechLM papers.
3.1 Speech Tokenizer
@@
(section 3), color=brightlavender, fill=brightlavender!15, text width=2.2cm, text=black [Speech Tokenizer
@@
(section 3.1), color=brightlavender, fill=brightlavender!15, text width=2.1cm, text=black [ Contrastive Predictive Coding Oord et al. (2018), Wav2vec 2.0 Baevski et al. (2020), vq-wav2vec Baevski et al. (2019), w2v-bert Chung et al. (2021), HuBERT Hsu et al. (2021), SpeechTokenizer Zhang et al. (2023b), SoundStream Zeghidour et al. (2021), Google USM Zhang et al. (2023c), WavLM Chen et al. (2022a) , color=brightlavender, fill=brightlavender!15, text width=8cm, text=black] ], [Language Model
@@
(section 3.2), color=brightlavender, fill=brightlavender!15, text width=2.1cm, text=black [ Transformer Vaswani (2017), LLaMA Touvron et al. (2023a), LLaMA 2 Touvron et al. (2023b), LLaMA 3 Dubey et al. (2024a), OPT Zhang et al. (2022b) , color=brightlavender, fill=brightlavender!15, text width=8cm, text=black], ], [Vocoder
@@
(section 3.3), color=brightlavender, fill=brightlavender!15, text width=2.1cm, text=black [ WaveNet Oord et al. (2016), Tacotron 2 Shen et al. (2018), WaveGlow Prenger et al. (2019), HiFi-GAN Kong et al. (2020a), Token-based HiFi-GAN Polyak et al. (2021), EnCodec Défossez et al. (2022) , color=brightlavender, fill=brightlavender!15, text width=8cm, text=black], ], ] [Training Recipes
@@
(table 2), color=lightgreen, fill=lightgreen!15, text width=2.2cm, text=black [Features Modeled
@@
(section 4.1), color=lightgreen, fill=lightgreen!15, text width=2.2cm, text=black [Discrete Features
@@
(section 4.1.1), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ Semantic Tokens:
@@
GSLM Lakhotia et al. (2021), TWIST Hassid et al. (2024), SpeechGPT Zhang et al. (2023a), AudioPaLM Rubenstein et al. (2023) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], [ Paralinguistic Token:
@@
pGSLM Kharitonov et al. (2021), SPIRIT-LM Nguyen et al. (2024) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], [ Acoustic Token:
@@
VioLA Wang et al. (2024), Li et al. (2024) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], [ Mixed Token:
@@
Moshi Défossez et al. (2024), SpeechGPT-Gen Zhang et al. (2024) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ], [Continuous Features
@@
(section 4.1.2), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ Spectron Nachmani et al. (2023), tGSLM Algayres et al. (2023), Mini-Omni Xie and Wu (2024), LauraGPT Du et al. (2023) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] ], [Training Stages
@@
(section 4.2), color=lightgreen, fill=lightgreen!15, text width=2.2cm, text=black [LM Pre-Training
@@
(section 4.2), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ Cold Initialization:
@@
GSLM Lakhotia et al. (2021), SUTLM Chou et al. (2023), pGSLM Kharitonov et al. (2021), LSLM Ma et al. (2024), VioLA Wang et al. (2024) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], [ Continued Pre-Training:
@@
TWIST Hassid et al. (2024), AudioPaLM Rubenstein et al. (2023), SPIRIT-LM Nguyen et al. (2024), AudioChatLlama Fathullah et al. (2024), Spectron Nachmani et al. (2023) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black] ], [LM Instruction-Tuning
@@
(section 4.2), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ SpeechGPT Zhang et al. (2023a), SpeechGPT-Gen Zhang et al. (2024), COSMIC Pan et al. (2023), Llama-Omni Fang et al. (2024) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] ], [Speech Generation Paradigm
@@
(section 4.3), color=lightgreen, fill=lightgreen!15, text width=2.2cm, text=black [Traditional Generation Paradigm
@@
(section 4.3), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ TWIST Hassid et al. (2024), SPIRIT-LM Nguyen et al. (2024), AudioPaLM Rubenstein et al. (2023), SpeechGPT Zhang et al. (2023a) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ], [Real-Time Interaction
@@
(section 4.3), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ dGSLM Nguyen et al. (2023b), LSLM Ma et al. (2024), VITA Fu et al. (2024), Moshi Défossez et al. (2024) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] [Silence Mode
@@
(section 4.3), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ VITA Fu et al. (2024) , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] ] ] ]
@@
Speech tokenizer is the first component in SpeechLMs, which encodes continuous audio signals (waveforms) into latent representations and then converts the latent representations into discrete tokens (or sometimes called speech units). This conversion allows the audio input to be effectively processed by a language model for tasks such as speech recognition or synthesis. Speech tokenizer aims to capture essential features of the audio while reducing its dimensionality, facilitating the subsequent modeling and analysis of speech patterns. In this section, we categorize speech tokenizers based on their focus on modeling different aspects of the raw audio.
3.1.1 Semantic Understanding Objective

### v2

@@ Abstract
Large Language Models (LLMs) have recently garnered significant attention, primarily for their capabilities in text-based interactions. However, natural human interaction often relies on speech, necessitating a shift towards voice-based models. A straightforward approach to achieve this involves a pipeline of “Automatic Speech Recognition (ASR) + LLM + Text-to-Speech (TTS)”, where input speech is transcribed to text, processed by an LLM, and then converted back to speech. Despite being straightforward, this method suffers from inherent limitations, such as information loss during modality conversion, significant latency due to the complex pipeline, and error accumulation across the three stages. To address these issues, Speech Language Models (SpeechLMs)—end-to-end models that generate speech without converting from text—have emerged as a promising alternative. This survey paper provides the first comprehensive overview of recent methodologies for constructing SpeechLMs, detailing the key components of their architecture and the various training recipes integral to their development. Additionally, we systematically survey the various capabilities of SpeechLMs, categorize their evaluation metrics, and discuss the challenges and future research directions in this rapidly evolving field. [^1] Github: https://github.com/dreamtheater123/Awesome-SpeechLM-Survey
Index Terms:
Speech Language Models, Speech Interaction, Large Language Models.
I Introduction
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities in generating text and performing a wide array of natural language processing tasks [1, 2, 3], serving as powerful foundation models for AI-driven language understanding and generation. Their success has also spurred numerous applications in various other domains, yet the reliance solely on text-based modalities presents a significant limitation. This leads to the development of speech-based generative models, which allow to interact with humans more naturally and intuitively. The inclusion of speech not only facilitates real-time voice interactions but also enriches communication by combining both text and speech information [4, 5].
@@
Given the extensive mutual information between text and speech, it is natural to modify existing LLMs to enable speech interaction capabilities. A straightforward approach is to adopt an “Automatic Speech Recognition (ASR) + LLM + Text-to-Speech (TTS)” framework (Figure 1a) [6, 7]. In this setup, the user’s spoken input is first processed by the ASR module, which converts it into text. The LLM then generates a text response based on this transcription. Finally, the TTS module transforms the text response back into speech, which is played back to the user. However, this naive solution mainly suffers from the following three problems.
  1. 1.
    Information loss. Speech signals not only contain semantic information (i.e., the meaning of the speech) but also paralinguistic information (e.g., pitch, timbre, tonality, etc.). Putting a text-only LLM in the middle will cause the complete loss of paralinguistic information in the input speech [8, 9]. Recognizing paralinguistic features allows for a more engaging and immersive interaction, as the model can respond with greater context. Moreover, it enables the model to interpret a user’s intent more accurately in certain situations, as the meaning of speech can vary depending on tone.
  2. 2.
    Significant latency. The sequential operation of ASR, LLM, and TTS leads to considerable delays due to the inherently complex structures and pipelines of these modules [10, 9, 11]. For instance, ASR often includes an additional text generator [12, 13], while TTS typically relies on a text tokenizer, both of which increase computational demands. Moreover, implementing advanced decoding techniques, such as beam search for ASR text decoding, can further contribute to delays [12, 13]. Consequently, developing an end-to-end speech generation model can help significantly minimize latency.
  3. 3.
    Cumulative error. A staged approach like this can easily lead to cumulative errors throughout the pipeline, particularly in the ASR-LLM stage [14, 15]. Specifically, transcription errors that occur when converting speech to text in the ASR module can negatively impact the language generation performance of the LLM. Additionally, the TTS performance can deteriorate significantly if the LLM generates text that cannot be synthesized. Consequently, developing a unified architecture plays a crucial role in reducing accumulated errors.
TABLE I: Notations used in this survey.
Notation Description
a A speech audio waveform
$\hat{\textbf{a}}$ Model-reconstructed speech audio waveform
t A text span
M A multi-modal sequence containing speech and/or text
$\theta$ Model parameters
$f_{E}(\cdot)$ Speech encoder
$d(\cdot)$ Speech quantizer
$LM$ Language model
v Encoded speech representations
s Speech tokens
m Multi-modal tokens (speech and/or text)
$V$ Vocabulary of a language model
$E$ Embedding matrix of a language model
De Transformer decoder blocks
$Vo$ Vocoder
$G$ Generator in a Generative Adversarial Network
$D$ Discriminator in a Generative Adversarial Network
$ms$ Mel-spectrogram of a speech audio waveform
$F_{0}$ Fundamental frequency
@@
The limitations of the naive ASR + LLM + TTS framework have led to the development of Speech Language Models (SpeechLMs, Figure 1b). Unlike the naive framework, SpeechLMs directly encode speech waveforms into tokens or representations, capturing essential features and information from audio (section III-A). Although individual speech tokens may not carry word-level semantic meaning, they capture the semantic information of speech utterances and retain valuable paralinguistic information, which prevents the information loss. SpeechLMs then model these tokens autoregressively, without solely relying on text input, which allows them to use the additional paralinguistic information to generate more expressive and nuanced speech (section III-B). Finally, the generated tokens are synthesized back to speech (section III-C). This integrated approach eliminates the need to chain together three separate modules, significantly reducing latency. Additionally, by working directly with the encoded speech tokens, SpeechLMs effectively mitigate the cumulative errors, as their training is integrated with the speech encoding, whereas the training of LLMs (language modeling) is completely independent of the ASR (speech recognition) module in the naive framework.
@@
Speech Language Models (SpeechLMs) have the capability to go beyond simple conversational tasks and address more intricate and diverse applications. First, SpeechLMs can capture speaker-specific information and emotional nuances (Figure 2), which allows them to distinguish between different speakers during a conversation and to comprehend and generate speech imbued with specific emotional tones. Such advancements are crucial for applications in areas like personalized assistants, emotion-aware systems, and more nuanced human-computer interaction scenarios. Second, SpeechLMs can be designed to enable real-time voice interaction, where the model can be interrupted by humans or choose to speak while the user is still speaking, mimicking the dynamics of human conversations more naturally. Furthermore, because SpeechLMs are trained directly on speech data, they have the potential to facilitate communication in rare languages where spoken content is more prevalent than written material.
@@
Contributions. In this survey, we present the first comprehensive overview of recent endeavors in constructing SpeechLMs. We explore the various components that constitute their architecture (section III) and the training recipes (table III) involved in their development. we aim to elucidate the current state of the field by analyzing these models from the above perspectives. Additionally, we survey the downstream applications of SpeechLMs (section V), classify metrics to evaluate SpeechLMs (section VI), discuss the challenges encountered in this rapidly evolving area, and outline promising future research directions that could drive further advancements in SpeechLM technology (section VII). Our contributions are summarized as follows:
@@ II Problem Formulation
In this section, we provide a formal definition of Speech Language Models. A Speech Language Model (SpeechLM) is an autoregressive foundation model that processes and generates speech end-to-end, utilizing contextual understanding for coherent sequence generation. This capability allows it to perform a variety of tasks through speech-based interactions. Although SpeechLMs are required to perform end-to-end speech interactions, they can also incorporate text, enabling cross-modal functionalities such as speech-in-text-out and vice versa. We note that the concept of SpeechLM is in contrast to traditional text-based language models, such as LLM, where the only modality being processed within the model is text. Therefore, to avoid confusion, we call those text-based language models TextLMs throughout this survey.
We offer a unified framework in which SpeechLMs can process and generate speech data, text data, or even interleaved speech and text data. Specifically, a speech audio waveform $\textbf{a}=(a_{1},a_{2},\ldots,a_{Q})$ consists of a sequence of audio samples $a_{i}\in\mathbb{R}$ of length $Q$, where $1\leq q\leq Q$. Similarly, a text span $\textbf{t}=(t_{1},t_{2},\ldots,t_{K})$ consists of a sequence of text tokens $t_{j}$ (word, subword, character, etc.) of length $K$. Let $\textbf{M}=(M_{1},M_{2},\ldots,M_{N})$ denote a multimodal sequence of length $N$, where each element $M_{i}\in\{a_{i},t_{j}\}$. We define $\textbf{M}^{\text{in}}=(M_{1}^{\text{in}},M_{2}^{\text{in}},\ldots,M_{N_{\text{in}}}^{\text{in}})$ as the input multimodal sequence and $\textbf{M}^{\text{out}}=(M_{1}^{\text{out}},M_{2}^{\text{out}},\ldots,M_{N_{\text{out}}}^{\text{out}})$ as the output multimodal sequence, where $N_{\text{in}}\geq 0$ and $N_{\text{out}}\geq 0$. Then, A SpeechLM parameterized by $\theta$ can then be represented as:
@@
III Components in SpeechLM
There are three main components within a SpeechLM, namely speech tokenizer, language model, and token-to-speech synthesizer (vocoder), as illustrated in Figure 1. The fundamental reason for such a three-staged design pattern is to use the language modeling architecture (e.g., decoder-only transformer) to model speech autoregressively in the format of audio waveforms. Since both the input and output of a language model are tokens, additional modules need to be attached to the language model to handle the I/O format. Specifically, the speech tokenizer first transforms continuous audio waveforms into tokens or representations to serve as input to the language model, then the language model performs the next-token prediction based on the input speech tokens. Finally, the vocoder transforms the tokens outputted by the language model back into audio waveforms. We note that our focus here is on how the three components are grouped together to form a SpeechLM rather than a comprehensive overview of each component. Therefore, for speech tokenizer and vocoder, we mainly summarize the methods used in existing SpeechLMs. Table II summarizes the popular choices of the three components in various SpeechLM papers.
III-A Speech Tokenizer
@@
(section III), color=brightlavender, fill=brightlavender!15, text width=2.2cm, text=black [Speech Tokenizer
@@
(section III-A), color=brightlavender, fill=brightlavender!15, text width=2.1cm, text=black [ Contrastive Predictive Coding [16], Wav2vec 2.0 [17], vq-wav2vec [18], w2v-bert [19], HuBERT [20], SpeechTokenizer [21], Mimi [9], SoundStream [22], Google USM [23], WavLM [24] , color=brightlavender, fill=brightlavender!15, text width=8cm, text=black] ], [Language Model
@@
(section III-B), color=brightlavender, fill=brightlavender!15, text width=2.1cm, text=black [ Transformer [25], LLaMA [26], LLaMA 2 [27], LLaMA 3 [2], OPT [3], Qwen2 [28], GLM [29], Mixtral [30] , color=brightlavender, fill=brightlavender!15, text width=8cm, text=black], ], [Vocoder
@@
(section III-C), color=brightlavender, fill=brightlavender!15, text width=2.1cm, text=black [ WaveNet [31], Tacotron 2 [32], WaveGlow [33], HiFi-GAN [34], Token-based HiFi-GAN [35], EnCodec [36] , color=brightlavender, fill=brightlavender!15, text width=8cm, text=black], ], ] [Training Recipes
@@
(table III), color=lightgreen, fill=lightgreen!15, text width=2.2cm, text=black [Features Modeled
@@
(section IV-A), color=lightgreen, fill=lightgreen!15, text width=2.2cm, text=black [Discrete Features
@@
(section IV-A1), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ Semantic Tokens:
@@
GSLM [37], TWIST [38], SpeechGPT [8], AudioPaLM [39], OmniFlatten [40], SLAM-Omni [41] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], [ Paralinguistic Token:
@@
pGSLM [42], SPIRIT-LM [5] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], [ Acoustic Token:
@@
VioLA [43], Li et al. [44], Parrot [45] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], [ Mixed Token:
@@
Moshi [9], SpeechGPT-Gen [46] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ], [Continuous Features
@@
(section IV-A2), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ Spectron [47], tGSLM [48], Mini-Omni [10], LauraGPT [49], SLAM-Omni [41] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] ], [Training Stages
@@
(section IV-B), color=lightgreen, fill=lightgreen!15, text width=2.2cm, text=black [LM Pre-Training
@@
(section IV-B), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ Cold Initialization:
@@
GSLM [37], SUTLM [50], pGSLM [42], LSLM [51], VioLA [43] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], [ Continued Pre-Training:
@@
TWIST [38], AudioPaLM [39], SPIRIT-LM [5], AudioChatLlama [14], Spectron [47], SLAM-Omni [41], OmniFlatten [40], Mini-Omni [10], Mini-Omni 2 [52], Freeze-Omni [53], Moshi [9], Parrot [45] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black] ], [LM Instruction-Tuning
@@
(section IV-B), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ SpeechGPT [8], SpeechGPT-Gen [46], COSMIC [54], Llama-Omni [11], Moshi [9] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] ], [Speech Generation Paradigm
@@
(section IV-C), color=lightgreen, fill=lightgreen!15, text width=2.2cm, text=black [Traditional Generation Paradigm
@@
(section IV-C), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ TWIST [38], SPIRIT-LM [5], AudioPaLM [39], SpeechGPT [8] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ], [Real-Time Interaction
@@
(section IV-C), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ dGSLM [4], Parrot [45], LSLM [51], VITA [55], Moshi [9], Mini-Omni 2 [52], OmniFlatten [40], SALMONN-omni [56] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] [Silence Mode
@@
(section IV-C), color=lightgreen, fill=lightgreen!15, text width=2.1cm, text=black [ VITA [55] , color=lightgreen, fill=lightgreen!15, text width=5.3cm, text=black], ] ] ] ]
@@
Speech tokenizer is the first component in SpeechLMs, which encodes continuous audio signals (waveforms) into tokens. Speech tokenizer aims to capture essential features of the audio while reducing its dimensionality and allows the audio input to be effectively processed by a language model for autoregressive generation. Speech tokenizer operates by encoding the audio segment by segment, producing two possible types of tokens (features): discrete tokens and continuous tokens. Discrete tokens (Section IV-A1) use a specific index to represent each speech segment, while continuous tokens (Section IV-A2) use an embedding to represent the segment [^2] For simplicity, we unify both types of representations as ”tokens”. Both token types can be utilized as input for a language model in autoregressive modeling. In this section, we categorize speech tokenizers based on their focus on modeling different aspects of the raw audio.
III-A1 Semantic Understanding Objective